Wilcox1Evee:

Baxter is thinking about buying a car for $18,500. The table below shows the projected value of two different cars for three years. Number of years 1 2 3 Car 1 (value in dollars) 17,390 16,346.60 15,365.80 Car 2 (value in dollars) 17,500 16,500 15,500 Part A: What type of function, linear or exponential, can be used to describe the value of each of the cars after a fixed number of years? Explain your answer. (2 points) Part B: Write one function for each car to describe the value of the car f(x), in dollars, after x years. (4 points) Part C: Baxter wants to purchase a car that would have the greatest value in 9 years. Will there be any significant difference in the value of either car after 9 years? Explain your answer, and show the value of each car after 9 years. (4 points)

Angelwolf:

Part a: is a Linear function Part B: Y=1000×X + 19463 Y = -1076.7×X + 19602 Part c: Car #1 will be worth 628$ more than car 2 after 10 years

Wilcox1Evee:

Thank you! Do you think you could explain your answers for Part A and Part C?

Angelwolf:

Part C-Car #1 Y = -1,000×X + 19,463, Y = 10×-1000 + 19,463 = 9,463 Car #2 Y = X×-1076.7 + 19602, Y = -1076.7×10 + 19,602 = 8,835 Part A-You have to form a plot of the data, the linear function is the values of both the cars after the fixed number.

Wilcox1Evee:

OK Thank you so much!!! Is there anything i can help you with?

Angelwolf:

Nah i'm good thanks for offering tho just glad i could help :)

Wilcox1Evee:

OK:))